An engine raises 100kg of water through a height of 60m in 20s what is the power of the engine? Take g as 10ms
1
Expert's answer
2020-09-24T11:07:27-0400

By the definition of the power we have:


"P = \dfrac{W}{t} = \dfrac{Fh}{t} = \dfrac{mgh}{t},""P = \dfrac{100 \ kg \cdot 10 \ \dfrac{m}{s^2} \cdot 60 \ m}{20 \ s} = 3 \cdot 10^3 \ W = 3 \ kW."

Answer:

"P =3 \cdot 10^3 \ W = 3 \ kW."